# Fault Diagnosis in New Wind Turbines using Knowledge from Existing Turbines by Generative Domain Adaptation

### v1

@@ 5.2 Detailed analysis of a domain pair example
We present a detailed comparison of the fault diagnosis performance across all methods for one specific domain pair with the source domain WT “WT07” and target domain WT “WT05” with 1 month of available training data. The performance gains for this domain pair represent approximately the average for this scarcity scenario.
@@ 5.2.1 NBM trained on representative data
To obtain our ground truth anomaly scores for the target WT test set, we trained an NBM on the full target training data, which contains measurements recorded over a span of more than 2 years (Appendix D). We visualize the resulting anomaly scores, i.e., the NBM reconstruction errors, for the target WT test set in Figure 6. The scores generally exceed the threshold during time frames when incidents were logged, yet not for all incidents (e.g., in mid-October 2023). As the incidents are unspecified, we are unable to categorize this as a missed fault or correct fault diagnosis. During normal operation, the scores tend to remain below the threshold, although certain short periods exhibit elevated scores (e.g.,in the middle of January 2024), which we cannot determine as false positives or identified but unlogged abnormal states. However, we only focus on comparing these scores to the ones obtained trained with scarce data.
@@ 5.2.2 NBM trained on scarce data
An NBM was then trained on only 1 month of target training data preceding the identical test set. The anomaly scores of this model are visualized in Figure 7. When comparing the threshold similarity, this model achieved an F1-score of 63.9% (Table A4). Notably, numerous anomaly scores falsely exceed the threshold compared to their counterpart in the ground truth NBM (e.g., January 2024), while scores during incidents largely match the ones from the representative NBM. As autoencoders learn normal behavior from the training set, this could indicate that normal operational states were missing in the limited training set, causing high reconstruction errors at test time and therefore falsely elevated anomaly scores. Overall, a significant decrease in fault diagnosis performance can be observed when only scarce training data is available.
@@ 5.2.3 Fine-tuning the source domain NBM
As a domain adaptation benchmark, an NBM was first trained on the abundant source training data, i.e., the full training set of WT07, and then fine-tuned using 1 month of target domain training data. Results are shown in Figure 8. This model achieved an increase in F1-score to 68.8% (+4.9 percentage points) compared to the scarce NBM. In particular, we notice a decrease in falsely elevated scores (for instance, October 2023) while anomaly scores generally remain elevated when they are in the ground truth. Our results suggest that fine-tuning is capable of adjusting information embedded in the source domain NBM to the target domain, resulting in unseen normal data missing in the target WT to be still considered normal, enabling more reliable fault diagnosis when training data is lacking.
@@ 6 Conclusion
This study investigated the application of generative domain adaptation, specifically domain mapping, to address the challenge of limited training data when using normal behavior models in wind turbine fault diagnosis. NBMs require substantial normal operation data, which when missing can result in unreliable diagnosis models. To address this challenge, we proposed a novel domain adaptation approach that leverages data from a WT of a different wind farm for which abundant training data is available. Our domain mapping technique transforms SCADA data from a data scarce target domain WT to resemble data from a different WT with representative training data and thus a reliable NBM. This mapping allows for the use of the source WT’s NBM for fault diagnosis in the target WT despite limited data. We validated our method by conducting experiments using 28 different WT pairs across varying degrees of data scarcity (1 to 8 weeks). Our findings demonstrate the superior performance of generative domain adaptation compared to training models on limited data and conventional fine-tuning, particularly when faced with severe data scarcity. Our results highlight the significant potential of domain mapping in achieving earlier and more reliable WT fault diagnosis models when only scarce training data is available, for instance in newly installed wind farms.
@@
Our study has several limitations which provide opportunities for future research. First, further exploration of time-series-based model architectures and frameworks, consistency losses, and generalizable hyperparameter optimization would be useful to further investigate the potential of domain mapping for WT fault diagnosis. In our study, hyperparameter tuning was evaluated on the test set of one selected domain pair using 1 month of target training data. It is possible that certain types of WT pairs exhibiting varying domain shifts may require differently weighted loss constraints and models. More generalizable optimization techniques could be investigated, e.g., based on a quantification of the domain shift. Our domain mapping network furthermore exhibits a significantly higher complexity and a computationally heavier and longer training procedure compared to fine-tuning, representing a consideration for practical applications. While we validated our approach on a very comprehensive dataset comprising real operational data from numerous WTs from different wind farms, our experiments were limited to WTs of the same manufacturer and a selection of SCADA features. Further large-scale experiments with more versatile wind farms and WT types, as well as different SCADA variables and systems, will be valuable to assess the generalizability of our results to other manufacturers. Lastly, evaluating our method under various types of faults (if logs are available) could provide insights in the limitations of mapping types of anomalies and lead to further improvements to the domain adaptation approach.
@@
Our exploratory work suggests a promising potential for further domain mapping applications. For WT fault diagnosis, this includes investigating the use of different architectures and adjusted techniques to establish domain mapping as a novel and effective alternative to fine-tuning. The proposed technique moreover highlights a potential for applications beyond WT systems. Exploring its applicability to diagnosis tasks under data scarcity of other areas, such as for photovoltaic systems, and generally anomaly detection tasks, is an interesting area of future research to explore domain mapping for more accurate and reliable models when faced with limited training data.
@@ Acknowledgements
All our autoencoder-based NBMs follow the model specifications outlined in Table A1, based on a model architecture search on a randomly chosen WT. Each NBM was trained using the RAdam [RAdam] optimizer with a learning rate of 0.003, minimizing the mean squared reconstruction error between model input and autoencoder reconstruction. The models were trained using a batch size of 128 SCADA samples. Training was stopped once the reconstruction error on the validation set stopped improved for several data scarcity-dependent epochs (250 for 2 weeks, 25 for full target data).
@@
Block 1 Conv1d (32 filters, kernel size 7, stride 1, Mish [misraMishSelfRegularized2019]) x 2
@@
GroupNorm [Wu_2018_ECCV] (1 group, 32 channels)
@@
The used architecture for the domain mapping discriminators is outlined in Table A2. For the generators, we used a residual temporal convolutional network (TCN) approach [bai_empirical_2018]. Our non-causal architecture consists of several residual TCN blocks, kernel size, dilation, and normalization, depicted in Figure A1. The full generator is outlined in Table A3.
@@
Figure A1: Structure of our TCN residual block, performing non-causal 1D convolutions with kernel size $k$, dilation $d$, and with $c$ channels. The final group normalization is optional when Norm is set. Based on [bai_empirical_2018].
@@
We follow the GAN-QP formulation to train our generators and discriminators. Let $G$ be a generator, $T$ a discriminator, $x_{r}$ a real sample and $x_{f}$ a fake sample (in our case, a mapped sample). The GAN-QP loss is defined as [suGANQPNovelGAN2018]:
@@
Our domain mapping network was trained with a batch size of 128 using an Adam optimizer ($\beta_{1}=0.5,\beta_{2}=0.999$, learning rate = $0.0002$) for the generators and discriminators each. Early stopping was implemented to stop training at an optimal state and to prevent overfitting. We use a reconstruction error obtained from the source NBM, by mapping the target validation data to the source domain. As only normal data is used for training and validation, we expect the reconstruction error of mapped target-to-source data to be low, representing realistic source WT data familiar to the source NBM. A rising reconstruction error could indicate overfitting on training data and unrealistic sample generation, as the mapped validation data starts to resemble the source domain less. It should be noted that this value cannot be used for model selection and tuning, as it contains no information regarding the content preservation (e.g., mapping all target samples to the same normal source domain sample would result in a very low score). Training was stopped after 1000 batch iterations with no improvement of the early stopping score. Moreover, at evaluation time, we used an Exponential Moving Average (EMA) of the training weights for the generators, which improved training stability and performance (as investigated in [emagan]).

### v2

@@ 5.2 Detailed analysis of a domain pair example
We present a detailed comparison of the fault detection performance across all methods for one specific domain pair with the source domain WT “WT07” and target domain WT “WT05” with 1 month of available training data. The performance gains for this domain pair represent approximately the average for this scarcity scenario.
@@ 5.2.1 NBM trained on representative data
To obtain our ground truth anomaly scores for the target WT test set, we trained an NBM on the full target training data, which contains measurements recorded over a span of more than 2 years (Appendix D). We visualize the resulting anomaly scores, i.e., the NBM reconstruction errors, for the target WT test set in Figure 6. The scores generally exceed the threshold during time frames when incidents were logged, yet not for all incidents (e.g., in mid-October 2023). As the incidents are unspecified, we are unable to categorize this as a missed fault or correct fault detection. During normal operation, the scores tend to remain below the threshold, although certain short periods exhibit elevated scores (e.g.,in the middle of January 2024), which we cannot determine as false positives or identified but unlogged abnormal states. However, we only focus on comparing these scores to the ones obtained trained with scarce data.
@@ 5.2.2 NBM trained on scarce data
An NBM was then trained on only 1 month of target training data preceding the identical test set. The anomaly scores of this model are visualized in Figure 7. When comparing the threshold similarity, this model achieved an F1-score of 63.9% (Table A4). Notably, numerous anomaly scores falsely exceed the threshold compared to their counterpart in the ground truth NBM (e.g., January 2024), while scores during incidents largely match the ones from the representative NBM. As autoencoders learn normal behavior from the training set, this could indicate that normal operational states were missing in the limited training set, causing high reconstruction errors at test time and therefore falsely elevated anomaly scores. Overall, a significant decrease in fault detection performance can be observed when only scarce training data is available.
@@ 5.2.3 Fine-tuning the source domain NBM
As a domain adaptation benchmark, an NBM was first trained on the abundant source training data, i.e., the full training set of WT07, and then fine-tuned using 1 month of target domain training data. Results are shown in Figure 8. This model achieved an increase in F1-score to 68.8% (+4.9 percentage points) compared to the scarce NBM. In particular, we notice a decrease in falsely elevated scores (for instance, October 2023) while anomaly scores generally remain elevated when they are in the ground truth. Our results suggest that fine-tuning is capable of adjusting information embedded in the source domain NBM to the target domain, resulting in unseen normal data missing in the target WT to be still considered normal, enabling more reliable fault detection when training data is lacking.
@@ 6 Conclusion
This study investigated the application of generative domain adaptation, specifically domain mapping, to address the challenge of limited training data when using normal behavior models in wind turbine fault detection. NBMs require substantial normal operation data, which when missing can result in unreliable models. To address this challenge, we proposed a novel domain adaptation approach that leverages data from a data-rich source domain WT of a different wind farm. Our domain mapping technique transforms SCADA data from a data-scarce target domain WT to resemble data from a different WT with representative training data. This mapping allows for the use of the source WT’s NBM for fault detection in the target WT despite limited data. We validated our method by conducting experiments using 28 different WT pairs across varying degrees of data scarcity (1 to 8 weeks). Our findings demonstrate the superior performance of generative domain adaptation compared to training models on limited data and conventional fine-tuning, particularly when faced with severe data scarcity. Our results highlight the significant potential of domain mapping in achieving earlier and more reliable WT fault detection models when only scarce training data is available, for instance in newly installed wind farms.
@@
Our study has several limitations which provide opportunities for future research. First, further exploration of time-series-based model architectures and frameworks, consistency losses, and generalizable hyperparameter optimization would be useful to further investigate the potential of domain mapping for WT fault detection. In our study, hyperparameter tuning was evaluated on the test set of one selected domain pair using 1 month of target training data. It is possible that certain types of WT pairs exhibiting varying domain shifts may require differently weighted loss constraints and models, for which more generalizable optimization techniques could be investigated. Moreover, performance differences between WT pairs indicate a potential benefit in appropriately selecting suitable source-target WTs, which suggests needed future investigations to find appropriate selection criteria, e.g., based on the of the domain shift. Our domain mapping network furthermore exhibits a significantly higher complexity and a computationally heavier and longer training procedure compared to fine-tuning, which should be considered for practical applications. While we validated our approach on a very comprehensive dataset comprising real operational data from numerous WTs from different wind farms, our experiments were limited to WTs of the same manufacturer and a selection of SCADA features. Further large-scale experiments with more versatile wind farms and WT types, as well as different SCADA variables and systems, will be valuable to assess the generalizability of our results to other manufacturers. Lastly, evaluating our method when logs are available of various types of faults could provide insights into the limitations of mapping specific types of faults and lead to further improvements.
@@
Our exploratory work suggests a promising potential for further domain mapping applications. For WT fault detection, this includes investigating the use of different architectures and adjusted techniques to establish domain mapping as a novel and effective alternative to fine-tuning. The proposed technique moreover highlights a potential for applications beyond WT systems. Exploring its applicability to unsupervised anomaly detection-based tasks under data scarcity of other areas, such as for photovoltaic systems, is an interesting area of future research to explore domain mapping for more accurate and reliable models when faced with limited training data.
@@ Acknowledgements
All our autoencoder-based NBMs follow the model specifications outlined in Table A1, based on a model architecture search on a randomly chosen WT. Each NBM was trained using the RAdam [43] optimizer with a learning rate of 0.003, minimizing the mean squared reconstruction error between model input and autoencoder reconstruction. The models were trained using a batch size of 128 SCADA samples. Training was stopped once the reconstruction error on the validation set stopped improved for several data scarcity-dependent epochs (250 for 2 weeks, 25 for full target data).
@@
Block 1 Conv1d (32 filters, kernel size 7, stride 1, Mish [44]) x 2
@@
GroupNorm [45] (1 group, 32 channels)
@@
The used architecture for the domain mapping discriminators is outlined in Table A2. For the generators, we used a residual temporal convolutional network (TCN) approach [46]. Our non-causal architecture consists of several residual TCN blocks, kernel size, dilation, and normalization, depicted in Figure A1. The full generator is outlined in Table A3.
@@
Figure A1: Structure of our TCN residual block, performing non-causal 1D convolutions with kernel size $k$, dilation $d$, and with $c$ channels. The final group normalization is optional when Norm is set. Based on [46].
@@
We follow the GAN-QP formulation to train our generators and discriminators. Let $G$ be a generator, $T$ a discriminator, $x_{r}$ a real sample and $x_{f}$ a fake sample (in our case, a mapped sample). The GAN-QP loss is defined as [41]:
@@
Our domain mapping network was trained with a batch size of 128 using an Adam optimizer ($\beta_{1}=0.5,\beta_{2}=0.999$, learning rate = $0.0002$) for the generators and discriminators each. Early stopping was implemented to stop training at an optimal state and to prevent overfitting. We use a reconstruction error obtained from the source NBM, by mapping the target validation data to the source domain. As only normal data is used for training and validation, we expect the reconstruction error of mapped target-to-source data to be low, representing realistic source WT data familiar to the source NBM. A rising reconstruction error could indicate overfitting on training data and unrealistic sample generation, as the mapped validation data starts to resemble the source domain less. It should be noted that this value cannot be used for model selection and tuning, as it contains no information regarding the content preservation (e.g., mapping all target samples to the same normal source domain sample would result in a very low score). Training was stopped after 1000 batch iterations with no improvement of the early stopping score. Moreover, at evaluation time, we used an Exponential Moving Average (EMA) of the training weights for the generators, which improved training stability and performance (as investigated in [47]).